OpenStudy (anonymous):

Solve: 3(x + 4) greater than or equal to x - 2

OpenStudy (anonymous):

Is it x >= -3?

jimthompson5910 (jim_thompson5910):

\[\large 3(x + 4) \ge x-2\] \[\large 3x + 12 \ge x-2\] \[\large 3x \ge x-2-12\] \[\large 3x-x \ge -2-12\] \[\large 2x \ge -14\] \[\large x \ge \frac{-14}{2}\] \[\large x \ge -7\]

jimthompson5910 (jim_thompson5910):

not sure how you got x >= -3

OpenStudy (anonymous):

Oh I know!! When I was distributing 3, I forgot to multiply it with 4, so instead of 12 I just kept the 4!! THANK YOU FOR HELPING ME!!!!! ^_^

jimthompson5910 (jim_thompson5910):

ah I gotcha, that gets a lot of people all the time it might help to make a table like this